in a right triangle the legs are 3 rational 3 find the length of the hypothenuse
Answer by RAY100(1637)

Right triangle therefore pythagorous applies,,,,c^2= a^2 +b^2
.
c^2 = (3sqrt3)^2 + (3sqrt3)^2
.
c^2 = 9*3 +9*3 = 27 +27 =54 =9*6
.
c=3sqrt6
.
check,,(3sqrt6)^2 = (3sqrt3)^2 + (3sqrt3)^2
7.348^2 =5.196^2 + 5.196^2
54 =27 +27 =54,,,,,ok